If f(x)=(2/3)x+2 find f(x-3) and write it in terms of f(x)
Answer by fractalier(6550)

f(x-3) = (2/3)(x-3) + 2 = (2x-6)/3 + 6/3 = 2x/3 = (2/3)x = f(x) - 2